I need help solving this problem
-5≤2x+5 and 2x+5<9
The solution is {X|██≤X<██}

Answer by Theo(13342)

-5≤2x+5 and 2x+5<9
This is equivalent to:

-5 <= 2x+5 < 9

whatever you do to one side of the equation, you have to do to all the other sides.

subtract 5 from all sides of the equation to get:

-5 -5 <= 2x + 5 - 5 < 9 - 5

this becomes:

-10 <= 2x < 4

divide all sides of this equation by 2 to get:

-5 <= x < 2

test in the original equation to see if the original equation holds true.

take several values of x within the limits and outside of the limits.

try x = -6, -5, 0, 1, 2

-6 and 2 are outside the limits. -5, 0, 1 are within the limits.

when x = -6, equation of -5 <= 2x+5 < 9 becomes:
-5 <= -12 + 5 < 9 which becomes:
-5 <= -7 < 9 which is NOT true because -5 is NOT <= -7
this is ok because -6 was outside the limits.

when x = -5, equation of -5 <= 2x+5 < 9 becomes:
-5 <= -10 + 5 < 9 which becomes:
-5 <= -5 < 9 which IS true.
this is ok because -5 was within the limits.

when x = 0, equation of -5 <= 2x+5 < 9 becomes:
-5 <= 0 + 5 <= 9 which becomes:
-5 <= 0 <= 9 which IS true.
this is ok because 0 was within the limits.

when x = 1, equation of -5 <= 2x+5 < 9 becomes:
-5 <= 2 + 5 < 9 which becomes:
-5 <= 7 < 9 which IS true.
this is ok because 1 was within the limits.

when x = 2, equation of -5 <= 2x+5 < 9 becomes:
-5 <= 4 + 5 < 9 which becomes:
-5 <= 9 < 9 which is NOT true because 9 is not < 9.
this is ok because 2 was outside the limits.

equation looks good and your answer is:

-5 <= x < 2

which is the same as:

-5 <= x and x < 2.

this is also the same as:

x >= -5 and x < 2

you could also have done each of the inequalities separately and then combined them.

your equation was:

-5 <= 2x + 5 and 2x+5 < 9

solve -5 <= 2x + 5 to get -5 <= x

solve 2x + 5 < 9 to get x < 2

-5 <= x and x < 2 is equivalent to -5 <= x < 9.